OpenStudy (anonymous):

Solve. However, I am more interested in knowing why my method does not work: Set u=arcsinx du=1/sqroot(1-x^2)dx integral of (sin(u)+u) du.

OpenStudy (anonymous):

Your method is fine, but your result is incorrect. You made a mistake with computing \(\int u~du\)

OpenStudy (anonymous):

So where you have \(\int (\sin u+u)~du\), you should have \(-\cos u+\dfrac{u^2}{2}+C\)

OpenStudy (anonymous):

I corrected it. :) But, when I graph the result from my teacher and my result, the slopes are different.

OpenStudy (anonymous):

Yes, that is what I got.

OpenStudy (anonymous):

Right, so back-substituting yields \(-\cos(\arcsin x)+\dfrac{\arcsin^2x}{2}+C\), which is the same as your teacher's result.

OpenStudy (anonymous):

You are a genius.

OpenStudy (anonymous):

I get it all now. THANKS. :)

OpenStudy (anonymous):

yw!